Find the area of the curve r^2 = 3-4cos^2∅, symmetrical to origin and Ox.


1
Expert's answer
2021-05-07T12:32:16-0400

Given, the curve r2 = 3-4cos2∅, symmetrical to origin.

Area of region above the origin=

"=\frac{1}{2}\int_{\frac{\pi}{6}}^{\frac{5\pi}{6}} r^2 d\varnothing\newline
=\frac{1}{2}\int_{\frac{\pi}{6}}^{\frac{5\pi}{6}} 3-4cos^2\varnothing d\varnothing\newline
=\frac{1}{2}\int_{\frac{\pi}{6}}^{\frac{5\pi}{6}} 3-4\frac{1-cos2\varnothing}{2} d\varnothing\newline
=\frac{1}{2}\int_{\frac{\pi}{6}}^{\frac{5\pi}{6}} 1+cos2\varnothing d\varnothing\newline
=\frac{1}{2}[ \varnothing + \frac{sin2\varnothing}{2} ]_{\frac{\pi}{6}}^{\frac{5\pi}{6}}\newline
=\frac{1}{2}[ \frac{5\pi}{6} -\frac{\sqrt3}{4} -{\frac{\pi}{6}}-\frac{\sqrt3}{4}]\newline
=\frac{1}{2}[ \frac{2\pi}{3} -\frac{\sqrt3}{2}]\newline
=0.614 \text{square units}"

Total area=2×0.614=1.228 square units.